Two angles are supplementary. The larger angle is 166 larger than the smaller angle. Find the measure of both angles.
Answer #1

Let x and y be the two angles of supplementary.

Supplementary angles add to 180°

Thus, x + y = 180°

Given x = 166° + y when x > y

Equation becomes (166° + y) + y = 180°

166° + 2y = 180°

2y = 180° - 166°

2y = 14° implies y = 14° / 2

Thus, y = 7°

x = 166 + y = 166° + 7 °

x = 173°

Hence, the two supplementary angles are 173° and 7°